OpenStudy (anonymous):

explain how to Find the amplitude of y = 3 sin x.

OpenStudy (anonymous):

y = A * sin(bx) A = amplitude what is it?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

you got it see how easy it was

OpenStudy (anonymous):

Describe the transformations required to obtain the graph of the function f(x) from the graph of the function g(x). f(x) = 9 cos x; g(x) = cos x

OpenStudy (anonymous):

what do you multiply g(x) to get f(x)?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

A.Horizontal stretch by a factor of 9 B.Vertical stretch by a factor of 9 C.Horizontal shrink by a factor of 1over9 D.Vertical shrink by a factor of 1over9

OpenStudy (anonymous):

B.Vertical stretch by a factor of 9

OpenStudy (anonymous):

you can graph them and see the change

OpenStudy (anonymous):

how do you graph y=2/3sin1/2x

OpenStudy (anonymous):

2/3 is the amplitude correct? plug in values for x y=2/3sin1/2x x=0; y = 0 can you try it then i will go over it with you

OpenStudy (anonymous):

this is a differnt question it says find thet graph that goes with y= 2 --- sin 1 3 ---- 2 x

OpenStudy (anonymous):

is that the question?\[\frac{ 2 }{ 3 }*\sin(\frac{ 1 }{ 2x })\]

OpenStudy (anonymous):

sara you there?

OpenStudy (anonymous):

its not 2x its 1/2 x

OpenStudy (anonymous):

1 --- x 2

OpenStudy (anonymous):

ok amplitude is 2/3 period is \[0 \le x \le \frac{ 2*\Pi }{ 1/2 }\]

OpenStudy (anonymous):

plug in 0 what do you get \[\frac{ 2 }{ 3 }*\sin{\frac{ 1 }{ 2 }*0}\]

OpenStudy (anonymous):

+ you have the amplitude; which graph shows the right amp?

OpenStudy (anonymous):

how do i know which graph shows 2/3

OpenStudy (anonymous):

all graphs has marked where 1 is